OpenStudy (anonymous):

PLEASE HELP-I have 2 questions! 1-Know the Trigonometric ratios and how they are used. 2-Be sure to know their reciprocal functions

OpenStudy (unklerhaukus):

those arn't questions

OpenStudy (anonymous):

well, i kinda wrote it in my own form. but i need to know what trigonometric ratios are and how they are used. And, i need to know the reciprocal functions

OpenStudy (anonymous):

i dont understand what reciprocal functions are and the 2st problem as well.

OpenStudy (anonymous):

1st*

OpenStudy (unklerhaukus):

well have you herd if the sine function \[\sin(\theta)=\frac{opp}{hypo}\] the sine of an angle is a measure of the ratio of 'the opposite side' to 'the hypotenuse' of a right angled triangle